The table below shows the price of a Grande latte from Starbucks Coffee in five different countries. The table also shows the official exchange rate between each currency and the dollar. Calculate the implied exchange rate in units of foreign currency per dollar, using the latte prices 1. Countiyv Price Latte Implied Official Exchange Exchange Rate Rate United States S3.55 Switzerland 6.90 franc 194 francs 0.96 franc r $ r S Sweden 37.46 krona 8.34 krona r S China 27 yuan 6.27 yuan r S India 150 rupee 62.37 rupee er S Compared to the U.S., prices in Switzerland are 102.1 percent higher (circle one) If you took $50,000 to Switzerland it would feel like you had S 24,742.27 purchasing power parity adjustment. after Compared to the U.S., prices in India are you took S50,000 to India it would feel like you had S parity adjustment. percent higher/lower (circle one). If after purchasing power

Need help, correct answers are shown for the first one. All I need to know is how to arrive at those answers, please show work.
Answer #1
Country Price Latte Implied Exchange Rate Official Exchange Rate
United States 3.55
Switzerland 6.9 1.94 0.96
Sweden 37.46 5.43 8.34
China 27 0.72 6.27
India 150 5.56 62.37

Implied exchange rate = Price of Latte in country 'X'/ Price of Latte in US

Compared to the US prices in India are = (5.56 - 62.37)/62.37 x 100 = - 91.1% (lower). (= (implied ER - official ER)/ official ER x 100)

If you took $ 50000 to India it would feel like you had = 50000 x 62.37/5.56 = $ 560881.3 ( = (50000 x official ER/implied ER)